# Practical hybrid decoding scheme for parity-encoded spin systems

### v3

@@ III.2 Readout decoding of the SLHZ model
If $\hat{\boldsymbol{r}}^{(n)}\rightarrow\hat{\boldsymbol{z}}$, or in other words, $\hat{\boldsymbol{e}}^{(n)}\rightarrow\hat{\boldsymbol{1}}_{K\times K}$ at $n=n_{0}$, where $\hat{\boldsymbol{1}}_{K\times K}$ is $K\times K$ matrix whose entries are all one (all-one matrix), we say that the decoding succeeded after $n_{0}$ iteration of the BF decoding. In the case of success, it follows that $s_{ijk}^{(3)}(\hat{\boldsymbol{r}}^{(n)})=1$ for a set of possible $\left\{i,j,k\right\}$ at $n=n_{0}$. Our algorithm can be viewed as an iterative process that gradually increases the number of syndromes with unit values by flipping spins according to the majority vote of the syndromes associated with each spin.
@@ IV Experimental demonstration
We demonstrate that the BF algorithm is valid for eliminating readout errors in the SLHZ model arising not only from i.i.d. noise but also from non-i.i.d. noise.
IV.1 In the case of i.i.d. noise
@@
The performance of the BF algorithm was investigated under the assumption of i.i.d. noise, the simplest model of noisy readout in QA. Pastawski and Preskill previously studied this model using the BP algorithm [20]. Note that the BF decoding in Eq.(36) depends only on the weight-3 syndromes, which treat all variables (VN and CN) symmetrically [20]. Furthermore, if we note that the LDPC codes are linear codes and assume the AWGN channel is symmetric, i.e. $P(y_{i}|z_{i})=P(-y_{i}|-z_{i})$ holds, the probability of successful decoding is independent of the input code-state $\boldsymbol{z}$ [31]. In this case, without loss of generality, the performance of decoding can be analyzed using the assumption that all-one code-state $\boldsymbol{z}=\left(+1,\ldots,+1\right)$ has been transmitted [31, 32]. In physics, this assumption corresponds to choosing the ferromagnetic gauge and identifying the current decision $\boldsymbol{r}$ with an error pattern $\boldsymbol{e}$. Thus, we assume the input is all-one code-state, that is, $\hat{\boldsymbol{z}}=\hat{\boldsymbol{1}}_{K\times K}$. All the demonstrations were performed using the Mathematica® Ver.14 platform on the Windows 11 operating system. We generated 5000 symmetric matrices $\hat{\boldsymbol{r}}=\hat{\boldsymbol{e}}$ with unit diagonal elements and other elements randomly assigned 1-1 with probability $\varepsilon<\tfrac{1}{2}$ and +1+1 otherwise. After $n=5$ iterations, it was checked whether the errors were eliminated. Figure 6(a) shows the performance of the BF algorithm, plotting the probability of decoding failure as a function of the number $K$ of logical spins ranging from $2$ to $40$ (associated SLHZ model consists of $\tbinom{K}{2}$ physical spins) for seven values of common bit error rate $\varepsilon$ $\left(=0.05,0.07,0.1,0.15,0.2,0.3,0.4\right)$. The MATLAB code implementing the described BF decoder and producing the figure is included as Supplemental Material [*[SeeSupplementalMaterialat][forMATLABcodeimplementationofdescribedBFdecoding.]matlabcode]. The failure probability falls steeply as $N$ increases if $\varepsilon$ is not too close to the threshold value $1/2$. A similar performance calculation was reproduced for the decoding using the BP algorithm after being iterated five times, as shown in Fig.6(b) [20]. Comparing these figures shows that the performance of the BF algorithm is comparable to that of the BP algorithm. Note that the BP algorithm updates all marginal probabilities $P\left(x_{i}|\boldsymbol{y}\right)$ associated with the $\tbinom{K}{2}$ spins sequentially by passing a real-valued message between the associated VNs and CNs per a single iteration. In contrast, in the BF algorithm, all the $\tbinom{K}{2}$ variables $r_{ij}$ in the matrix $\hat{\boldsymbol{r}}$ are updated in parallel per a single iteration. Thus, each spin variable was updated five times in these algorithms.
@@
Figure 6: Comparison of performance of (a) BF, (b) BP, and (c) MCMC decodings. Assuming that the error probability $\varepsilon$ is common to all physical spins, the average probabilities of decoding failure are plotted as a function of the number $K$ of logical spins (associated with $\tbinom{K}{2}$ physical spins) for seven values of $\varepsilon$. Each data point was obtained by averaging over 5000 error matrix realizations. The BF and BP algorithms were iterated five times for each realization, and the MCMC sampling was iterated $\tbinom{K}{2}$ times. We considered a tie in the majority voting to be a failure in the BF decoding, although we can resolve it by introducing a tie-breaking rule, such as coin tossing.
Fig.7 is an typical example for a successfully decoded result when $K=40$ and $\varepsilon=0.3$. In this figure, each entry of the error matrix $\hat{\boldsymbol{e}}$ is plotted after $n=1,\ldots,4$ rounds of iteration of the operation in Eq.(46). The blue pixels correspond to spins with error ($e_{ij}=-1$), the number of which gradually decreases as rounds of iteration are added. In this example, an error-free matrix was obtained after $n=3$ iterations. Such results were observed for more than $70$ % of the error matrices generated.
@@
Figure 7: A typical example of successfully decoded results by the BF decoding when $K=40$ and $\varepsilon=0.3$. The initial readout was generated according to the i.i.d. noise model. The estimated matrix $\hat{\boldsymbol{r}}$ is plotted from left to right in increasing order of iterations $n$. The blue pixels represent spins with error.
@@
In addition to the BF and BP algorithms, let us focus on the word MAP decoding introduced in Sec.III.1 and compare its performance with the above two algorithms. The word MAP decoding says that any code-state is the ground state of the Hamiltonian, as shown in Eq.(13),
$H^{code}(\hat{\boldsymbol{x}})\equiv\gamma\sum_{\{i,j,k\}}\frac{1-s_{ijk}^{(3)}(\hat{\boldsymbol{x}})}{2},$ (48)
@@
$\hat{\boldsymbol{z}}=\underset{\hat{\boldsymbol{x}}\in\{\pm 1\}^{K\times K}}{\arg\min}H^{code}(\hat{\boldsymbol{x}})=\hat{\boldsymbol{1}}_{K\times K}.$ (49)
@@
IV.2 In the case of non-i.i.d. noise
Next, we consider the problem of primary interest in our study: what is expected when applying the BF algorithm to the readout of the SLHZ-based QA device. As the development of QA devices is ongoing and classical simulation of a quantum-mechanical many-body system is computationally challenging, it is difficult to investigate the potential of our BF decoding algorithm when applied to the readouts of an SLHZ-based QA device, both in actual QA devices and in classical computer simulations. In this study, we simulate spin readout in the SLHZ system using data stochastically sampled by a classical MCMC sampler. This analysis enables us to investigate the tolerance of the SLHZ model to leakage errors arising from dynamical and thermal excitations during the sampling.
We used the following Hamiltonian to sample data from the SLHZ system stochastically:
$H^{code}(\hat{\boldsymbol{x}})\equiv-\beta\sum_{\{i,j\}}J_{ij}x_{ij}+\gamma\sum_{\{i,j,k,l\}}\frac{1-s_{ijkl}^{(4)}(\hat{\boldsymbol{x}})}{2},$ (50)
where $J_{ij}\in\mathbb{R}$ is a coupling constant which is identified with the channel observation $y_{ij}\in\mathbb{R}$ in the AWGN channel model. The first term in Eq. (50) corresponds to the distance metric between a codeword $\boldsymbol{x}$ and the received information $\boldsymbol{J}$, while the second term corresponds to the constraint satisfaction metric [35]. Parameters $\left\{\beta,\gamma\right\}$ are the Lagrange weights for these two terms.
@@
Here, we must draw attention to coupling constants $\{J_{ij}\}$. Recall that the primary purpose of this study is to propose and demonstrate a practical method for readout decoding of the SLHZ-based QA devices. In this context, $\{J_{ij}\}$ is given by the problem we want to tackle. For example, they are given by an externally defined COP. In other cases, we can consider minimization of the Hamiltonian (50) as a decoding of the observation $\boldsymbol{J}=(J_{12},\ldots,J_{K-1\,K})\in\mathbb{R}^{\tbinom{K}{2}}$ in the AWGN channel model. To avoid duplicative consideration in analyzing the potential of the BF algorithm for readout decoding of the SLHZ model, we disregarded the reliability information for each spin readout of the SLHZ model itself, even if it were available in the actual QA device. Thus, the following results can be considered valid if readout errors are solely due to leakage errors arising from dynamical and thermal excitations during the sampling process, while neglecting other errors, such as measurement errors. In the Appendix B, we present a supplementary result when i.i.d. noise is introduced during the readout stage. This demonstrates that the above simplified approach also exhibits reasonable robustness against i.i.d. readout errors.
@@
As a starting point, we provide typical examples demonstrating that both our BF and BP algorithms successfully eliminate errors in the readout of the SLHZ model generated by a classical MCMC sampler. In this example, we simulated readouts $\hat{\boldsymbol{r}}=\hat{\boldsymbol{e}}\circ\hat{\boldsymbol{z}}\in\{\pm 1\}^{K\times K}$ of the SLHZ system associated with a spin glass problem for $K=14$ ($K_{14}$) as a toy model. We generated a set of $12$ logical random instances on complete graphs $K_{14}$ with couplings $J_{ij}\in[-\tfrac{1}{4},\tfrac{1}{4}]$ chosen uniformly at random and all logical local fields set to zero, where the code-state $\hat{\boldsymbol{z}}$ for a given $\boldsymbol{J}$ was precomputed by brute force. The leftmost matrix plot in Fig.8 visualizes a typical example of the error matrix $\hat{\boldsymbol{e}}$ associated with a sampled readout $\hat{\boldsymbol{r}}$ of the SLHZ system. In this example, the error distribution was quite different from the expected for an i.i.d. noise model in Fig.7. Fig.8 shows the results where $\hat{\boldsymbol{e}}$ was decoded by the BF algorithm (upper) and the BP algorithm (lower) when the number of iterations was limited to 5 rounds. These results suggest that our BF algorithm, as well as the BP algorithm, can eliminate errors in the stochastically sampled readouts of the SLHZ system, which does not follow the i.i.d. noise model. Here, although we can not know the error rate for each spin because the correct ground state is generally unknown a priori, an initial error rate of 0.25 was assumed for every spin when running the BP algorithm.
Figure 9: Landscapes of the probability distribution for successful decoding. They are plotted as functions of the Lagrange weights $\{\beta,\gamma\}$ in Eq.(50). The left and right columns are the results of (a) MCMC decoding and (b) MCMC-BF hybrid decoding, respectively. The target states in the upper and lower rows of these figures are different: they are the correct code-state $\hat{\boldsymbol{z}}$ for the upper row and any code-states for the lower row.
Figure 10: Matrix plots representing the averaged error matrix $\langle\hat{\boldsymbol{e}}\rangle=\langle\hat{\boldsymbol{x}}\rangle\circ\hat{\boldsymbol{z}}$ for the sampled states by the MCMC sampling, where each entry reflects the marginal error probability in inferring the correct sign of the associated entry. Warm (cold) colored pixels indicate that there is likely no error (error) in the inference of the associated spin. The left and right plots correspond to $\langle\hat{\boldsymbol{e}}\rangle$ when $\{\hat{\boldsymbol{x}}\}$ was sampled under the parameter sets $A$ and $B$ in Fig.9, respectively.
@@
The performance of the two methods differs significantly. The columns (a) and (b) in Fig.9 indicate landscapes of the success probability for (a) the MCMC decoding and (b) the MCMC-BF hybrid decoding, respectively. In these figures, the top two figures are the success probability for finding the error-free code-state $\hat{\boldsymbol{z}}$, and the bottom two figures are the probability for finding any code-state $\hat{\boldsymbol{x}}=\boldsymbol{X}^{T}\boldsymbol{X}$, where $\boldsymbol{X}\in\{\pm 1\}^{K}$ is any logical state. It is very important to note that the sizes of the sample $\{\hat{\boldsymbol{x}}\}$ for the MCMC decoding and the decoded sample $\{\hat{\boldsymbol{r}}\}$ for the MCMC-BF hybrid decoding differ significantly; they were $1200\tbinom{K}{2}$ in the MCMC decoding and $4\tbinom{K}{2}$ in the MCMC-BF hybrid decoding. This indicates that the number of iterations required for MCMC sampling was 300 times smaller for the MCMC-BF hybrid decoding than for the MCMC decoding. Therefore, the MCMC-BF hybrid decoding provides a better trade-off between error performance and decoding complexity if the decoding complexity of the BF decoding is negligible. We will revisit the validity of this assumption later.
@@
Let us note two arrows $A$ and $B$ shown in Fig.9. They correspond to the sets of Lagrange weights $A=\{\beta_{A},\gamma_{A}\}$ and $B=\{\beta_{B},\gamma_{B}\}$ for which the success probability for decoding is maximized in the MCMC decoding (A) and the MCMC-BF hybrid decoding (B), respectively. The figure clearly indicates that the parameter sets suitable for use differ between the respective decoding methods. The left and right in Fig. 10 are the matrix plots showing the average error matrix $\langle\hat{\boldsymbol{e}}\rangle=\langle\hat{\boldsymbol{x}}\rangle\circ\hat{\boldsymbol{z}}$ over the set of samples $\{\hat{\boldsymbol{x}}\}$ obtained from the MCMC sampler under the parameter sets $A$ and $B$, respectively. The marginal probability $P_{c}$ that $\hat{x}_{ij}$ gives the correct value, i.e., $\hat{x}_{ij}=\hat{z}_{ij}$, is given by $P_{c}=\tfrac{1+\langle{\hat{e}_{ij}}\rangle}{2}$ . So, a negative entry in $\langle\hat{\boldsymbol{e}}\rangle$ indicates it is likely that there is an error in the corresponding entry in $\hat{\boldsymbol{x}}$. We can see that $\langle\hat{\boldsymbol{e}}\rangle$ depends on the chosen parameter sets. The average error matrix $\langle\hat{\boldsymbol{e}}\rangle$ in Fig.10(b) is consistent with the error matrix $\hat{\boldsymbol{e}}$ in Fig.8, because they were sampled with the same parameter set $B$. Note that it is highly likely to fail sampling a code-state by MCMC sampler for the set $B$ (see the lower left plot in Fig.9), while it is highly likely to succeed sampling a code-state by the MCMC sampler for the set $A$. This implies that the state $\hat{\boldsymbol{x}}$ sampled in the first stage of the MCMC-BF decoding is not a code-state, although the state $\hat{\boldsymbol{x}}$ sampled in the MCMC decoding is a code-state. This is quite reasonable because the subsequent BF decoding can correct errors only if $\hat{\boldsymbol{x}}$ sampled at the first stage is a leakage state in the MCMC-BF hybrid decoding. In contrast, the sampled state must be a code-state if it is the correct state in the MCMC decoding. This is why optimal parameter sets $\{\beta_{opt},\gamma_{opt}\}$ differ between the MCMC decoding and MCMC-BF hybrid decoding. It is important to note that this is a consequence of our BF decoding algorithm and is independent of the sampling algorithm in the first stage. Our results suggest that optimal Lagrange weights should be carefully chosen when applying our BF decoding for readouts of QA as they may differ from the optimal parameters for QA when used alone.
V Discussions

### v4

@@ III.2 Readout decoding of the SLHZ model
If $\hat{\textbf{r}}^{(n)}\rightarrow\hat{\textbf{z}}$, or in other words, $\hat{\textbf{e}}^{(n)}\rightarrow\hat{\textbf{1}}_{K\times K}$ at $n=n_{0}$, where $\hat{\boldsymbol{1}}_{K\times K}$ is $K\times K$ matrix the entries of which are all one (all-one matrix), we say that the decoding has succeeded after $n_{0}$ iterations of the BF decoding. In the case of success, it follows that $s_{ijk}^{(3)}(\hat{\textbf{r}}^{(n)})=1$ for a set of possible $\{i,j,k\}$ at $n=n_{0}$. Our algorithm can be viewed as an iterative process that gradually increases the number of syndromes with unit values by flipping spins according to the majority vote of the syndromes associated with each spin.
@@ IV EXPERIMENTAL DEMONSTRATION
We demonstrate that the BF algorithm is valid for eliminating readout errors in the SLHZ model arising not only from IID noise but also from non-IID noise.
IV.1 In the case of IID noise
@@
The performance of the BF algorithm has been investigated under the assumption of IID noise, the simplest model of noisy readout in QA. Pastawski and Preskill have previously studied this model using the BP algorithm [20]. Note that the BF decoding in Eq.(36) depends only on the weight-3 syndromes, which treat all variables (VN and CN) symmetrically [20]. Furthermore, if we note that the LDPC codes are linear codes and assume that the AWGN channel is symmetric, i.e. $P(y_{i}|z_{i})=P(-y_{i}|-z_{i})$ holds, the probability of successful decoding is independent of the input code state z [30]. In this case, without loss of generality, the performance of decoding can be analyzed using the assumption that the all-one code state $\textbf{z}=\left(+1,\ldots,+1\right)$ has been transmitted [30, 31]. In physics, this assumption corresponds to choosing the ferromagnetic gauge and identifying the current decision r with an error pattern e. Thus, we assume the input is all-one code-state, that is, $\hat{\textbf{z}}=\hat{\textbf{1}}_{K\times K}$. All the demonstrations have been performed using the Mathematica® Ver.14 platform on the Windows 11 operating system. We have generated 5000 symmetric matrices $\hat{\textbf{r}}=\hat{\textbf{e}}$ with unit diagonal elements and other elements randomly assigned 1-1 with probability $\varepsilon<\tfrac{1}{2}$ and +1+1 otherwise. After $n=5$ iterations, we have checked whether the errors have been eliminated. In Fig. 6(a), we show the performance of the BF algorithm, plotting the probability of decoding failure as a function of the number $K$ of logical spins ranging from $2$ to $40$ (the associated SLHZ model consists of $\tbinom{K}{2}$ physical spins) for seven values of common bit error rate $\varepsilon$ $\left(=0.05,0.07,0.1,0.15,0.2,0.3,0.4\right)$. The MATLAB code implementing the described BF decoder and producing the figure is included as Supplemental Material [*[SeeSupplementalMaterialat][forMATLABcodeimplementationofdescribedBFdecoding.]matlabcode]. The failure probability falls steeply as $N$ increases if $\varepsilon$ is not too close to the threshold value $1/2$. A similar performance calculation has been reproduced for the decoding using the BP algorithm after being iterated 5 times, as shown in Fig.6(b) [20]. Comparing these figures shows that the performance of the BF algorithm is comparable to that of the BP algorithm. Note that the BP algorithm updates all marginal probabilities $P\left(x_{i}|\textbf{y}\right)$ associated with the $\tbinom{K}{2}$ spins sequentially by passing a real-valued message between the associated VNs and CNs per a single iteration. In contrast, in the BF algorithm, all the $\tbinom{K}{2}$ variables $r_{ij}$ in the matrix $\hat{\textbf{r}}$ are updated in parallel per a single iteration. Thus, each spin variable was updated five times in these algorithms. Fig.7 is an typical example for a successfully decoded result when $K=40$ and $\varepsilon=0.3$. In this figure, each entry of the error matrix $\hat{\textbf{e}}$ is plotted after $n=1,\ldots,4$ rounds of iteration of the operation in Eq.(46). The blue pixels correspond to spins with error ($e_{ij}=-1$), the number of which gradually decreases as rounds of iteration are added. In this example, an error-free matrix has been obtained after $n=3$ iterations. Such results have been observed for more than $70$ % of the error matrices generated.
@@
Figure 6: A comparison of the performance of the (a) bit-flipping (BF), (b) belief-propagation (BP), and (c) Markov-chain Monte Carlo (MCMC) decodings. Assuming that the error probability $\varepsilon$ is common to all physical spins, the average probabilities of decoding failure are plotted as a function of the number $K$ of logical spins (associated with $\tbinom{K}{2}$ physical spins) for seven values of $\varepsilon$. Each data point has been obtained by averaging over 5000 error-matrix realizations. The BF and BP algorithms have been iterated 5 times for each realization and the MCMC sampling has been iterated $\tbinom{K}{2}$ times. We have considered a tie in the majority voting to be a failure in the BF decoding, although we can resolve it by introducing a tie-breaking rule, such as coin tossing.
@@
Figure 7: A typical example of successfully decoded results by the BF decoding when $K=40$ and $\varepsilon=0.3$. The initial readout has been generated according to the IID-noise model. The estimated matrix $\hat{\textbf{r}}$ is plotted from left to right in increasing order of iterations $n$. The blue pixels represent spins with error.
@@
In addition to the BF and BP algorithms, let us focus on the word-MAP decoding introduced in Sec.III.1 and compare its performance with the above two algorithms. The word-MAP decoding says that any code state is the ground state of the Hamiltonian, as shown in Eq.(13),
$H^{code}(\hat{\textbf{x}})\equiv\gamma\sum_{\{i,j,k\}}\frac{1-s_{ijk}^{(3)}(\hat{\textbf{x}})}{2},$ (48)
@@
$\hat{\textbf{z}}=\underset{\hat{\textbf{x}}\in\{\pm 1\}^{K\times K}}{\arg\min}H^{code}(\hat{\textbf{x}})=\hat{\boldsymbol{1}}_{K\times K}.$ (49)
@@
IV.2 In the case of non-IID noise
Next, we consider the problem of primary interest in our study: what is expected when applying the BF algorithm to the readout of the SLHZ-based QA device. As the development of QA devices is ongoing and classical simulation of a quantum mechanical many-body system is computationally challenging, it is difficult to investigate the potential of our BF decoding algorithm when applied to the readouts of an SLHZ-based QA device, both in actual QA devices and in classical computer simulations. In this study, we simulate spin readout in the SLHZ system using data stochastically sampled by a classical MCMC sampler. This analysis enables us to investigate the tolerance of the SLHZ model to leakage errors arising from dynamical and thermal excitations during the sampling.
We have used the following Hamiltonian to sample data from the SLHZ system stochastically:
$H^{code}(\hat{\textbf{x}})\equiv-\beta\sum_{\{i,j\}}J_{ij}x_{ij}+\gamma\sum_{\{i,j,k,l\}}\frac{1-s_{ijkl}^{(4)}(\hat{\textbf{x}})}{2},$ (50)
where $J_{ij}\in\mathbb{R}$ is a coupling constant that is identified with the channel observation $y_{ij}\in\mathbb{R}$ in the AWGN channel model. The first term in Eq. (50) corresponds to the distance metric between a code word x and the received information J, while the second term corresponds to the constraint satisfaction metric [34]. Parameters $\left\{\beta,\gamma\right\}$ are the Lagrange weights for these two terms.
@@ IV.2 In the case of non-IID noise
Here, we must draw attention to coupling constants $\{J_{ij}\}$. Recall that the primary purpose of this study is to propose and demonstrate a practical method for readout decoding of the SLHZ-based QA devices. In this context, the $\{J_{ij}\}$ are given by the problem we want to tackle. For example, they are given by an externally defined COP. In other cases, we can consider minimization of the Hamiltonian in Eq,50 as a decoding of the observation $\textbf{J}=(J_{12},\ldots,J_{K-1\,K})\in\mathbb{R}^{\tbinom{K}{2}}$ in the AWGN channel model. To avoid duplicative consideration in analyzing the potential of the BF algorithm for readout decoding of the SLHZ model, we have disregarded the reliability information for each spin readout of the SLHZ model itself, even if it was available in the actual QA device. Thus, the following results can be considered valid if readout errors are solely due to leakage errors arising from dynamical and thermal excitations during the sampling process, while neglecting other errors, such as measurement errors. In Appendix B, we present a supplementary result when IID noise is introduced during the readout stage. This demonstrates that the above simplified approach also exhibits reasonable robustness against IID readout errors.
@@
As a starting point, we provide typical examples demonstrating that both our BF and BP algorithms successfully eliminate errors in the readout of the SLHZ model generated by a classical MCMC sampler. In this example, we have simulated readouts $\hat{\textbf{r}}=\hat{\textbf{e}}\circ\hat{\textbf{z}}\in\{\pm 1\}^{K\times K}$ of the SLHZ system associated with a spin-glass problem for $K=14$ ($K_{14}$) as a toy model. We have generated a set of $12$ logical random instances on complete graphs $K_{14}$ with couplings $J_{ij}\in[-\tfrac{1}{4},\tfrac{1}{4}]$ chosen uniformly at random and all logical local fields set to zero, where the code state $\hat{\textbf{z}}$ for a given J has been precomputed by brute force. The leftmost matrix plot in Fig.8 visualizes a typical example of the error matrix $\hat{\textbf{e}}$ associated with a sampled readout $\hat{\textbf{r}}$ of the SLHZ system. In this example, the error distribution is quite different from the expected one for an IID noise model in Fig.7. In Fig.8, we show the results where $\hat{\textbf{e}}$ has been decoded by the BF algorithm (upper) and the BP algorithm (lower) when the number of iterations was limited to five rounds. These results suggest that our BF algorithm, as well as the BP algorithm, can eliminate errors in the stochastically sampled readouts of the SLHZ system, which does not follow the IID-noise model. Here, although we cannot know the error rate for each spin because the correct ground state is generally unknown a priori, an initial error rate of 0.25 has been assumed for every spin when running the BP algorithm.
Figure 9: Landscapes of the probability distribution for successful decoding. They are plotted as functions of the Lagrange weights $\{\beta,\gamma\}$ in Eq.(50). The left and right columns are the results of (a) MCMC decoding and (b) MCMC-BF hybrid decoding, respectively. The target states in the upper and lower rows of these figures are different: they are the correct code state $\hat{\textbf{z}}$ for the upper row and any code states for the lower row.
Figure 10: Matrix plots representing the averaged error matrix $\langle\hat{\textbf{e}}\rangle=\langle\hat{\textbf{x}}\rangle\circ\hat{\textbf{z}}$ for the sampled states by the MCMC sampling, where each entry reflects the marginal error probability in inferring the correct sign of the associated entry. Warm- (cold-) colored pixels indicate that there is likely no error (error) in the inference of the associated spin. (a), (b) The plots correspond to $\langle\hat{\textbf{e}}\rangle$ when $\{\hat{\textbf{x}}\}$ has been sampled under the parameter sets (a) $A$ and (b) $B$ in Fig.9, respectively.
@@
The performance of the two methods differs significantly. Columns (a) and (b) in Fig.9 indicate landscapes of the success probability for (a) the MCMC decoding and (b) the MCMC-BF hybrid decoding, respectively. In Fig.9, the top row of figures shows the success probability of finding the error-free code state $\hat{\textbf{z}}$, and the bottom row of figures shows the probability for finding any code state $\hat{\textbf{x}}=\textbf{X}^{T}\textbf{X}$, where $\textbf{X}\in\{\pm 1\}^{K}$ is any logical state. It is very important to note that the sizes of the sample $\{\hat{\textbf{x}}\}$ for the MCMC decoding and the decoded sample $\{\hat{\textbf{r}}\}$ for the MCMC-BF hybrid decoding differ significantly; they are $1200\tbinom{K}{2}$ in the MCMC decoding and $4\tbinom{K}{2}$ in the MCMC-BF hybrid decoding. This indicates that the number of iterations required for MCMC sampling is 300 times smaller for the MCMC-BF hybrid decoding than for the MCMC decoding. Therefore, the MCMC-BF hybrid decoding provides a better trade-off between error performance and decoding complexity if the decoding complexity of the BF decoding is negligible. We will revisit the validity of this assumption later.
@@
Let us note two arrows $A$ and $B$ shown in Fig.9. They correspond to the sets of Lagrange weights $A=\{\beta_{A},\gamma_{A}\}$ and $B=\{\beta_{B},\gamma_{B}\}$ for which the success probability for decoding is maximized in the MCMC decoding (A) and the MCMC-BF hybrid decoding (B), respectively. The figure clearly indicates that the parameter sets suitable for use differ between the respective decoding methods. Figures 10(a) and 10(b) are matrix plots showing the average error matrix $\langle\hat{\textbf{e}}\rangle=\langle\hat{\textbf{x}}\rangle\circ\hat{\textbf{z}}$ over the set of samples $\{\hat{\textbf{x}}\}$ obtained from the MCMC sampler under the parameter sets $A$ and $B$, respectively. The marginal probability $P_{c}$ that $\hat{x}_{ij}$ gives the correct value, i.e., $\hat{x}_{ij}=\hat{z}_{ij}$, is given by $P_{c}=\tfrac{1+\langle{\hat{e}_{ij}}\rangle}{2}$ . So, a negative entry in $\langle\hat{\textbf{e}}\rangle$ indicates that it is likely that there is an error in the corresponding entry in $\hat{\textbf{x}}$. We can see that $\langle\hat{\textbf{e}}\rangle$ depends on the chosen parameter sets. The average error matrix $\langle\hat{\textbf{e}}\rangle$ in Fig.10(b) is consistent with the error matrix $\hat{\textbf{e}}$ in Fig.8, because they have been sampled with the same parameter set $B$. Note that it is highly likely to fail sampling a code state by the MCMC sampler for the set $B$ (see the lower left plot in Fig.9), while it is highly likely to succeed sampling a code state by the MCMC sampler for the set $A$. This implies that the state $\hat{\textbf{x}}$ sampled in the first stage of the MCMC-BF decoding is not a code state, although the state $\hat{\textbf{x}}$ sampled in the MCMC decoding is a code state. This is quite reasonable because the subsequent BF decoding can correct errors only if $\hat{\textbf{x}}$ sampled at the first stage is a leakage state in the MCMC-BF hybrid decoding. In contrast, the sampled state must be a code state if it is the correct state in the MCMC decoding. This is why the optimal parameter sets $\{\beta_{opt},\gamma_{opt}\}$ differ between the MCMC decoding and MCMC-BF hybrid decoding. It is important to note that this is a consequence of our BF decoding algorithm and thet it is independent of the sampling algorithm in the first stage. Our results suggest that optimal Lagrange weights should be carefully chosen when applying our BF decoding for readouts of QA, as they may differ from the optimal parameters for QA when used alone.
V DISCUSSION